please help solve and explain the following: 1. 2- (-3 + y) = 17
2. -6 X y/7 = -(72)
3. -(5 + /y/) = -3 /y/ =absolute value
4. -7y - 3(1/4y + 5) - 1/4(6y + 12) = 19

Answer by edjones(8007)

1. 2- (-3 + y) = 17
2+3-y=17
-y=12
y=-12
.
Ed